OpenStudy (anonymous):

If the average of a number N and its reciprocal is 1, what is the number N?

OpenStudy (lgbasallote):

the reciprocal of N is obviously \(\frac 1N\) so the average is 1 you say? \[\large \frac{N + \frac{1}{N}}{2} = 1\] \[\large N + \frac 1N = 2\] \[\large \frac{N^2+ 1}{N} = 2\] \[\large N^2 + 1 = 2N\] \[\large N^2 - 2N + 1 = 0\] solve for N

OpenStudy (anonymous):

so n=1

OpenStudy (lgbasallote):

yup